# On the minimax optimality of Flow Matching through the connection to kernel density estimation

### v1

@@ 3 Rate of convergence
The aim of this work is to evaluate how well Flow Matching performs depending on the number of observations $n.$ To this end, we have to control the distance between $\mathbb{P}^{*}$ and $\mathbb{P}^{\hat{\psi}_{1}(Z)}$ with flow $\hat{\psi}$ from (10). As as evaluation metric, we use the Wasserstein-$1$ metric.
On the normed vector space $(\mathbb{R}^{d},|\cdot|)$ the Wasserstein-$1$ distance between two probability distributions $\mathbb{P}$ and $\mathbb{Q}$ on $\mathbb{R}^{d}$ is defined as
@@
where $\Pi$ is the set of all distribution whose marginal distributions correspond to $\mathbb{P}$ and $\mathbb{Q}$ respectively. By [Villani, 2008, Theorem 5.10(i)], the following duality holds:
@@
On the space of probability measures with finite first moment, the Wasserstein-$1$ distance metrizes weak convergence. A comprehensive overview of the advantages of this distance over other measures that metrize weak convergence can be found in [Villani, 2008, p. 98 f.]. Due to its compatibility with dimension reduction settings such as the manifold hypothesis, the Wasserstein $1$ distance is a popular choice for the evaluation of generative models, see e.g. Schreuder et al. [2021], Liang [2021], Stéphanovitch et al. [2024] or Kunkel and Trabs [2025].
The following theorem provides a first comparison of the performance of $\mathbb{P}^{\hat{\psi}_{1}(Z)}$ and the empirical flow $\mathbb{P}^{\psi^{n}_{1}(Z)}$.
@@
Assume that all functions in $\mathcal{M}$ are Lipschitz continuous for fixed $t$ with Lipschitz constant $\Gamma_{t}.$ Then for any $\tilde{v}\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d},$ with $\tilde{v}\in\mathcal{M}$ and $v^{n}$ from (8)
@@
The first term in (15) depends on the choice of the conditional densities $p_{t}(\cdot|X_{i}^{*})$ and on the number of observations, but is independent of the approximation class $\mathcal{M}$. The error term $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})$ is a mixture of an approximation and a stochastic error and is determined by the choice of the flow $\psi^{n}_{1}$ or the distribution $p^{n}_{1}$, respectively. Due to the Flow Matching construction, the latter depends only on the conditional density path $p_{t}(\cdot|X_{i}^{*})$. Note that Lemma 2.3 demonstrates how $v_{t}(\cdot|X_{i}^{*})$ can be constructed based on a given $p_{t}(\cdot|X_{i}^{*})$. Using the construction from (11) to (13), $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})$ is the estimation error of a kernel density estimator.
The second term $\|v^{n}-\hat{v}\|_{\infty}$ depends on the set $\mathcal{M}$ and its ability to approximate $v^{n}$. For a sufficiently rich class $\mathcal{M}$ this second term will be negligible in our analysis.
The proof of Theorem 3.1 uses Grönwall’s Lemma, which leads to the factor $e^{\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t}$ in the second term in (15). This is standard in the analysis of ODEs, in context of flow-based generative models see Fukumizu et al. [2024] and Albergo and Vanden-Eijnden [2022]. Compared to the SDE setting of diffusion models, there is no such result like Girsanov’s theorem, which is typically used to study these models (following Chen et al. [2023c]) and leads to the avoidance of the exponential dependence on the Lipschitz constant. This greatly complicates the analysis of Flow Matching. Given this dependence, a Lipschitz regularization in the Flow Matching objective seems to be theoretically beneficial. This could be a possible avenue for future research.
@@
A bound for arbitrary Lipschitz $\tilde{v}_{t}$ cannot be better than $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})$. If $v^{n}_{t}$ is contained in $\mathcal{M},$ then the Picard-Lindelöf theorem yields $\hat{\psi}_{t}=\psi^{n}_{t}$ and hence $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{\psi_{1}}})=\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}}).$
Due to Theorem 3.1, we can bound bothnm error terms individually. We start with the error corresponding to the kernel density estimator in Section 3.1. Subsequently, we bound the second error term for approximation classes $\mathcal{M}$ consisting of feedforward neural networks in Section 3.2 to obtain our final result.
@@ 3.1 Rate of convergence of the KDE in Wasserstein distance
To bound $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi_{1}^{n}(Z)})$ in Theorem 3.1, we have to analyze the corresponding kernel density estimator in the Wasserstein-$1$ distance. For compact kernels, existing results for the $L_{1}$-error of the KDE can be exploited. However, this can only lead to the suboptimal convergence rate $n^{-\frac{\alpha}{2\alpha+d}}$. Indeed, the Lipschitz regularity of the test function in (14) can be exploited to obtain a faster rate. Additionally, to allow for latent distributions with unbounded support, especially the Gaussian distribution used by Lipman et al. [2023], we need to investigate kernels with unbounded support.
@@
Assume $d\geq 2,p^{*}\in B_{1,\infty}^{\alpha},\alpha\in(0,1]$ with $\|p^{*}\|_{B_{1,\infty}^{\alpha}}\leq M$ and $\operatorname{supp}(p^{*})$ bounded. Assume $K$, used in the Flow Matching setting via (13), is a nonnegative $d$-dimensional kernel such that
@@
$\mathbb{E}\big{[}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})\big{]}\leq C_{1}\sigma_{\min}^{1+\alpha}+\frac{C_{2}}{\sqrt{\sigma^{d-2}}}\frac{\log n}{\sqrt{n}}.$
For $\sigma_{\operatorname{min}}\sim(n/\log^{2}n)^{-\frac{1}{2\alpha+d}}$ we obtain $\mathbb{E}\big{[}\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})\big{]}=\mathcal{O}((n/\log^{2}n)^{-\frac{1+\alpha}{2\alpha+d}}).$
The convergence rate $(n/\log^{2}n)^{-\frac{1+\alpha}{2\alpha+d}}$ coincides up to the logarithmic factor with the lower bound by Niles-Weed and Berthet [2022, Theorem 3] and thus the above rate is minimax optimal up to the logarithm. Divol [2022] also considers rates of convergence for kernel density estimators. In case of $d\neq 2,$ he obtains the optimal rate for distributions bounded away from zero. However, the kernels he considers must be smooth radial functions with support bounded in $(0,1)^{d}$, [Divol, 2022, Condition A]. Hence, his result does not apply to the Gaussian kernel.
@@
    For kernels that to not satisfy the differentiability condition in Theorem 3.3, we show in Theorem A.1 a rate of convergence of $n^{-\frac{1+\alpha}{2\alpha+2+d}}.$ For nonzero $\alpha$ and $d$ big enough, this rate decays faster than $n^{-\frac{1}{d}},$ the rate of convergence of the empirical measure (Dudley [1969], Boissard and Gouic [2014]). This result replaces the assumption that $\operatorname{supp}(p^{*})$ is bounded by a moment assumption on $p^{*}.$
@@
    The second term in the proof of Theorem 3.3 is the expected value of $\mathsf{W}_{1}(K_{\sigma_{\min}}\ast\mathbb{P}_{n},K_{\sigma_{\min}}\ast\mathbb{P}^{*}).$ Results on the convergence of the smooth empirical measure have been discussed in case of the Gaussian kernel. Goldfeld et al. [2020, Proposition 1] obtain a rate in $\mathcal{O}\big{(}\frac{1}{\sqrt{n\sigma_{\min}^{d}}}\big{)}$ for subgaussian distributions $\mathbb{P}^{*}$, which coincides with the rate obtained in Theorem A.1. Weed [2018] looks at distributions on $[-1,1]$ and obtains a rate in $\mathcal{O}\big{(}\frac{1}{\sqrt{n}\sigma_{\min}^{d}}\big{)}.$ Hence the results above improves these results for distributions in $B^{\alpha}_{1,\infty}$ on bounded support.
@@ 3.2 Flow Matching with parameterized vector field
If the empirical vector field $v^{n}$ is contained in $\mathcal{M}$, then the second error term in Theorem 3.1 vanishes and the Flow Matching procedure inherits the optimal rate of convergence from Theorem 3.3. In practice, (9) is optimized over a class of neural networks to find a good approximate of $v^{n}$. We thus have to take this approximation error into account.
@@
    $\sigma_{t}(X_{i}^{*})=1-(1-\sigma_{\mathrm{min}})t\quad\text{and}\quad\mu_{t}(X_{i}^{*})=tX_{i}^{*}.$
@@
The choice of $\sigma_{t}$ and $\mu_{t}$ in 3.5 coincides with the setting of Lipman et al. [2023] in Example II. This is not a unique choice, for other choices used in the literature in similar settings, see [Tong et al., 2024, Table 1]. To avoid losing information, if $\mathbb{U}=\mathcal{N}(0,I_{d})$, then the choice of $\mu_{t}$ should satisfy $|\mu_{t}(X_{i}^{*})|\leq|X_{i}^{*}|$. $\sigma_{t}$ should map to $[0,1]$ to avoid unnecessarily increasing the Lipschitz constant $\Gamma_{t}$ of the vector field. Generalizing the linear choice of $\sigma_{t}$ to
$\sigma_{t}(X_{i}^{*})=(1-(1-\sigma_{\mathrm{min}}^{\frac{1}{k}}{})^{k}$
for a $k>0$ has only a negligible effect. The curvature cancels out in the construction of $v_{t}(\cdot|X_{i}^{*}),$ see Lemma 2.3.
In order to obtain quantitative results for the neural network approximation, we further need to specify the neural networks. We use fully connected feedforward ReLU networks, which enjoy nice approximation properties, see for example Yarotsky [2017], Gühring et al. [2020], Kohler and Langer [2021], Schmidt-Hieber [2020], Suzuki [2019].
Let us fix some notation: define the Rectified Linear Unit (ReLU) activation function, $\phi\colon\mathbb{R}\rightarrow\mathbb{R}$ $\phi(x)=\max(0,x)$. For a vector $v=\left(v_{1},\ldots,v_{p}\right)\in\mathbb{R}^{p}$, we define the shifted ReLU activation function $\phi_{v}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$ as
@@
Now we want to combine the approximation properties of ReLU networks with our above analysis to evaluate the performance of the Flow Matching mechanism from (10) where $\mathcal{M}=\mathsf{NN}(L,M,\Gamma)$ is a set of ReLU networks $v\colon[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d},(t,x)\mapsto v_{t}(x)$ with a fixed maximal number of layers $L$, at most $M$ nonzero weights and Lipschitz constant of $v_{t}$ of at most $\Gamma$ for any $t$. We obtain the following rate:
@@
Grant Assumption 3.5 and assume $\operatorname{supp}(p^{*})\subset[-1,1]^{d}$ and $p^{*}\in B^{\alpha}_{1,\infty}$ with $\|p^{*}\|_{B^{\alpha}_{1,\infty}}\leq M<\infty,\alpha\in(0,1]$. Set $\sigma_{\min}=n^{-\frac{1}{2\alpha+d}}.$ Then there are sequences $L_{n},M_{n},\Gamma_{n}\in\mathbb{N}$ such that for $n$ big enough
@@
where $\hat{\psi}$ is given by (10) with $\mathcal{M}=\mathsf{NN}(L_{n},M_{n},\Gamma_{n})$.
@@
    The Lipschitz constraint in Theorem 3.6 is typically not enforced in practice. However, we only require a bound on the Lipschitz constant of order $\Gamma_{n}=\frac{2d+1}{\sigma_{t}^{3}}+\frac{1}{2}$ which is only a mild restriction for $\sigma_{\min}\to 0$ as $n\to\infty$.
@@
    Compared to the literature on diffusion models and the first results on Flow Matching, we do not use early stopping times, but rather set $\sigma_{\min}>0$ according to $n.$ In terms of variance, both procedures are interchangeable: calculating $\sigma_{t_{*}}$ in [Fukumizu et al., 2024, Theorem 9] leads to exactly the same variance. However, early stopping leads to a bias induced by the mean function $\mu$, which is not $1$ at a time $t<1.$ This could be fixed using an adapted mean function, which is $1$ at the time of the early stopping, but this leads to much more complicated structures of $\mu.$
Theorem 3.6 shows that Flow Matching achieves minimax optimal rates (up to logarithmic factors) for certain classes of unknown densities, giving some justification of their the empirical success. There are several limitations and potential avenues for further research in the results of this work.
First of all the networks used for Theorem 3.6 are very large and do not correspond to the networks used in practice. While theoretical worst-case error bounds will always lead to networks that are not exactly small, the work of Yakovlev and Puchkin [2025] gives hope for improvement. Up to now it is an open question, whether a smaller network would come at the cost of a non-optimal rate. In addition, the size of the network and the overall result could benefit from the theoretical use of more sophisticated networks, like the U-net (Ronneberger et al. [2015]) construction in Lipman et al. [2023].

### v2

@@ 3 Rate of convergence
The aim of this work is to evaluate how well Flow Matching performs depending on the number of observations $n.$ To this end, we have to control the distance between $\mathbb{P}^{*}$ and $\mathbb{P}^{\hat{\psi}_{1}(Z)}$ with flow $\hat{\psi}$ from (11). As as evaluation metric, we use the Wasserstein $1$ metric.
On the normed vector space $(\mathbb{R}^{d},|\cdot|)$ the Wasserstein $1$ distance between two probability distributions $\mathbb{P}$ and $\mathbb{Q}$ on $\mathbb{R}^{d}$ is defined as
@@
where $\Pi$ is the set of all distributions whose marginal distributions correspond to $\mathbb{P}$ and $\mathbb{Q}$ respectively. By (Villani, 2008, Theorem 5.10(i)), the following duality holds:
@@
On the space of probability measures with finite first moments, the Wasserstein $1$ distance metrizes weak convergence. A comprehensive overview of the advantages of this distance over other measures that metrize weak convergence can be found in (Villani, 2008, p. 98 f.). Due to its compatibility with dimension reduction settings such as the manifold hypothesis, the Wasserstein $1$ distance is a popular choice for the evaluation of generative models, see e.g. Schreuder et al. (2021); Liang (2021); Stéphanovitch et al. (2024); Kunkel and Trabs (2025).
Another popular measure of dissimiliarity is the total variation distance,
$\operatorname{TV}(\mathbb{P},\mathbb{Q})\coloneqq\sup_{A\subset\mathcal{A}}|\mathbb{P}(A)-\mathbb{Q}(A)|,$
where $\mathcal{A}$ is the Borel-$\sigma$ algebra. This metric has frequently been used to evaluate diffusion models (Oko et al., 2023; Yakovlev and Puchkin, 2025). Example˜3.1 illustrates why the total variation distance is not suitable to evaluate Flow Matching models.
Example 3.1.
Let $d=1,Z\sim\mathcal{N}(0,1),\varepsilon>0$ and define the two vector fields $v^{(1)},v^{(2)}\colon[0,1]\times\mathbb{R}\rightarrow\mathbb{R}$
$v^{(1)}_{t}(x)\coloneqq 0,\quad v^{(2)}_{t}(x)\coloneqq\varepsilon\sin\Big(\frac{x}{\varepsilon}\Big),\quad\text{for all }t\in[0,1].$
Let $\psi^{(1)},\psi^{(2)}$ be the corresponding solutions to the ODE (3). Then
$\|v^{(1)}-v^{(2)}\|_{\infty}\overset{\varepsilon\rightarrow 0}{\rightarrow}0,\quad\mathsf{W}_{1}(\mathbb{P}^{\psi_{1}^{(1)}(Z)},\mathbb{P}^{\psi_{2}^{(1)}(Z)})\leq\varepsilon\overset{\varepsilon\rightarrow 0}{\rightarrow}0\quad\text{and}\quad\underset{\varepsilon\rightarrow 0}{\liminf}\operatorname{TV}(\mathbb{P}^{\psi_{1}^{(1)}(Z)},\mathbb{P}^{\psi_{2}^{(1)}(Z)})\geq c,$
for a constant $c>0$ .
The proof of Example˜3.1 shows that for convergence in total variation distance, one needs control over the gradient of $\psi$ with regard to $x$. Considering the change of variables theorem and the fact that the total variation distance operates on a density level, this is unsurprising. However, the gradient is not controlled by the Flow Matching objective. Assuming gradient control over the vector field in space can lead to different results, see Su et al. (2025).
The following theorem provides a first comparison of the performance of $\mathbb{P}^{\hat{\psi}_{1}(Z)}$ and the empirical flow $\mathbb{P}^{\psi^{n}_{1}(Z)}$ in the Wasserstein $1$ distance.
@@
Assume that all functions in $\mathcal{N}$ are Lipschitz continuous for fixed $t$ with Lipschitz constant $\Gamma_{t}.$ Then for any $\tilde{v}\colon[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d},$ with $\tilde{v}\in\mathcal{N}$ and $v^{n}$ from (9)
@@
Using the construction from (12) to (14), $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})$ is the estimation error of a kernel density estimator. The second term $\|v^{n}-\hat{v}\|_{\infty}$ depends on the set $\mathcal{N}$ and its ability to approximate $v^{n}$. For a sufficiently rich class $\mathcal{M}$ this second term will be negligible in our analysis. A corresponding error decomposition can be shown for the Wasserstein $2$ distance, too.
The proof of Theorem˜3.2 uses Grönwall’s Lemma, which leads to the factor $\exp(\int_{0}^{1}\Gamma_{t}\;\mathrm{d}t)$ in the second term in (16). This is standard in the analysis of ODEs, in context of flow-based generative models see Fukumizu et al. (2025) and Albergo and Vanden-Eijnden (2023). Compared to the SDE setting of diffusion models, there is no such result like Girsanov’s theorem, which is typically used to study these models, following Chen et al. (2023c), and circumvents the exponential dependence on the Lipschitz constant. This greatly complicates the analysis of Flow Matching. Given this dependence, a Lipschitz regularization in the Flow Matching objective seems to be theoretically beneficial. Imposing the assumption that the vector field defined in (7) is Lipschitz continuous implies an extensive restriction on the class of feasible $\mathbb{P}^{*}$, see Kunkel (2025).
@@
A bound for arbitrary Lipschitz $\tilde{v}_{t}$ cannot be better than $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})$. If $v^{n}_{t}$ is contained in $\mathcal{N},$ then the Picard-Lindelöf theorem yields $\hat{\psi}_{t}=\psi^{n}_{t}$ and hence $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\hat{\psi}_{1}})=\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}}).$
Due to Theorem˜3.2, we can bound both error terms individually. We start with the error corresponding to the kernel density estimator in Section˜3.1. Subsequently, we bound the second error term for approximation classes $\mathcal{M}$ consisting of feedforward neural networks in Section˜3.2 to obtain our final result.
@@ 3.1 Rate of convergence of the KDE in Wasserstein distance
To bound $\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi_{1}^{n}(Z)})$ in Theorem˜3.2, we have to analyze the corresponding kernel density estimator in the Wasserstein $1$ distance. For compact kernels, existing results for the $L_{1}$-error of the KDE can be exploited. However, this can only lead to the suboptimal convergence rate $n^{-\frac{\alpha}{2\alpha+d}}$. Indeed, the Lipschitz regularity of the test function in (15) can be exploited to obtain a faster rate. Additionally, to allow for latent distributions with unbounded support, especially the Gaussian distribution used by Lipman et al. (2023), we need to investigate kernels with unbounded support.
@@
Assume $d\geq 2,p^{*}\in B_{1,\infty}^{\alpha}\left(\Lambda,\mathbb{R}^{d}\right),\alpha\in(0,1],\Lambda>0$ and $\operatorname{supp}\left(p^{*}\right)$ bounded. Assume $K$, used in the Flow Matching setting via (14), is a nonnegative $d$-dimensional kernel such that
@@
$\mathbb{E}\big[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})\big]\leq C_{1}\sigma_{\min}^{1+\alpha}+\frac{C_{2}}{\sqrt{\sigma_{\min}^{d-2}}}\frac{\log n}{\sqrt{n}}.$
For $\sigma_{\min}\sim(n/\log^{2}n)^{-\frac{1}{2\alpha+d}}$ we obtain $\mathbb{E}\big[\mathsf{W}_{1}(\mathbb{P}^{*},\mathbb{P}^{\psi^{n}_{1}(Z)})\big]=\mathcal{O}((n/\log^{2}n)^{-\frac{1+\alpha}{2\alpha+d}}).$
The proof separates the bias induced by smoothing and the stochastic error after smoothing. Using the dual form (15) and the relationship between differentiation and convolution, the bound of the bias exploits both, the $\alpha$-smoothness of $\mathbb{P}^{*}$ and the Lipschitz-smoothness of test functions in (15). The the stochastic error is bounded by considering $(K_{\sigma_{\min}}\ast f)$ as the smoothed test-function and then taking advantage of a result by Schreuder (2020). Noting that the proof relies on the dual form of the Wasserstein $1$ distance, a bound for the Wasserstein $2$ distance would require a completly different approach.
The convergence rate $(n/\log^{2}n)^{-\frac{1+\alpha}{2\alpha+d}}$ coincides up to the logarithmic factor with the lower bound by Niles-Weed and Berthet (2022, Theorem 3) and thus the above rate is minimax optimal up to the logarithm. Divol (2022) also considers rates of convergence for kernel density estimators. In case of $d\neq 2,$ he obtains the optimal rate for distributions bounded away from zero. However, the kernels he considers must be smooth radial functions with support bounded in $(0,1)^{d}$, (Divol, 2022, Condition A). Hence, his result does not apply to the Gaussian kernel.
@@
    For kernels that to not satisfy the differentiability condition in Theorem˜3.4, we show in Theorem˜A.1 a rate of convergence of $n^{-\frac{1+\alpha}{2\alpha+2+d}}.$ For nonzero $\alpha$ and $d$ big enough, this rate decays faster than $n^{-\frac{1}{d}},$ the rate of convergence of the empirical measure (Dudley, 1969; Boissard and Gouic, 2014). This result replaces the assumption that $\operatorname{supp}(p^{*})$ is bounded by a moment assumption on $p^{*}.$
@@
    The second term in Theorem˜3.4 is an upper bound for the expected value of $\mathsf{W}_{1}(K_{\sigma_{\min}}\ast\mathbb{P}_{n},K_{\sigma_{\min}}\ast\mathbb{P}^{*}).$ Results on the convergence of the smooth empirical measure have been discussed in case of the Gaussian kernel. Goldfeld et al. (2020, Proposition 1) obtain a rate in $\mathcal{O}\big(\frac{1}{\sqrt{n\sigma_{\min}^{d}}}\big)$ for subgaussian distributions $\mathbb{P}^{*}$, which coincides with the rate obtained in Theorem˜A.1. Weed (2018) looks at distributions on [1,1][-1,1] and obtains a rate in $\mathcal{O}\big(\frac{1}{\sqrt{n}\sigma_{\min}^{d}}\big).$ Hence the results above improves these results for distributions in $B^{\alpha}_{1,\infty}$ on bounded support.
@@ 3.2 Flow Matching with parameterized vector field
If the empirical vector field $v^{n}$ is contained in $\mathcal{N}$, then the second error term in Theorem˜3.2 vanishes and the Flow Matching procedure inherits the optimal rate of convergence from Theorem˜3.4. In practice, (10) is optimized over a class of neural networks to find a good approximate of $v^{n}$. We thus have to take this approximation error into account.
@@
    $\sigma_{t}(X_{i}^{*})=1-(1-\sigma_{\min})t\quad\text{and}\quad\mu_{t}(X_{i}^{*})=tX_{i}^{*}.$
@@
The choice of $\sigma_{t}$ and $\mu_{t}$ in ˜3.6 coincides with the setting of Lipman et al. (2023) in Example II. This is not a unique choice, for other choices used in the literature in similar settings, see (Tong et al., 2024, Table 1). To avoid losing information, if $\mathbb{U}=\mathcal{N}(0,I_{d})$, then the choice of $\mu_{t}$ should satisfy $|\mu_{t}(X_{i}^{*})|\leq|X_{i}^{*}|$. $\sigma_{t}$ should map to $[0,1]$ to avoid unnecessarily increasing the Lipschitz constant $\Gamma_{t}$ of the vector field. Generalizing the linear choice of $\sigma_{t}$ to
$\sigma_{t}(X_{i}^{*})=(1-(1-\sigma_{\min}^{\frac{1}{k}}{})^{k}$
for a $k>0$ has only a negligible effect. The curvature cancels out in the construction of $v_{t}(\cdot|X_{i}^{*}),$ see Lemma˜2.3.
In order to obtain quantitative results for the neural network approximation, we further need to specify the neural networks. We use fully connected feedforward ReLU networks, which enjoy nice approximation properties, see for example Yarotsky (2017); Gühring et al. (2020); Kohler and Langer (2021); Schmidt-Hieber (2020); Suzuki (2019).
Let us fix some notation: define the Rectified Linear Unit (ReLU) activation function, $\phi\colon\mathbb{R}\rightarrow\mathbb{R},\phi(x)=\max(0,x)$. For a vector $v=\left(v_{1},\ldots,v_{p}\right)\in\mathbb{R}^{p}$, we define the shifted ReLU activation function $\phi_{v}:\mathbb{R}^{p}\rightarrow\mathbb{R}^{p}$ as
@@
Now we want to combine the approximation properties of ReLU networks with our above analysis to evaluate the performance of the Flow Matching mechanism from (11) where $\mathcal{N}=\mathsf{NN}(L,S,\Gamma)$ is a set of ReLU networks $v\colon[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d},(t,x)\mapsto v_{t}(x)$ with a fixed maximal number of layers $L$, at most $S$ nonzero weights and Lipschitz constant of $v_{t}$ of at most $\Gamma$ for any $t$. We obtain the following rate:
@@
Grant Assumption 3.6 and assume $p^{*}\in B^{\alpha}_{1,\infty}(\Lambda,[-1,1]^{d})$ for $\alpha\in(0,1],\Lambda>0$. Set $\sigma_{\min}=n^{-\frac{1}{2\alpha+d}}.$ Then we have for $n$ big enough
@@
where $\hat{\psi}$ is given by (11) with $\mathcal{N}=\mathsf{NN}(L_{n},S_{n},\Gamma_{n})$ where $L_{n}\geq O(n^{\ell})$, $S_{n}\geq O(\exp(n^{\ell}))$ and $\Gamma_{n}=O(n^{\frac{3}{2\alpha+d}})$ for some $\ell>0$.
The proof of Theorem˜3.7 combines Theorem˜3.2 with Theorem˜3.4, the tail behavior of the Gaussian distribution and an approximation result of Gühring et al. (2020). In a nutshell, we approximate the vector field $v^{n}$ on a growing but bounded set with an accuracy that compensates the exponential dependence on the Lipschitz constant. Given a Wasserstein $2$ version of Theorem˜3.4, the same approach could be used the extend the result to $\mathbb{E}[\mathsf{W}_{2}(\mathbb{P}^{*},\mathbb{P}^{\hat{\psi}_{1}(Z)})].$
@@
    The Lipschitz constraint in Theorem˜3.7 is typically not enforced in practice. However, we only require a bound on the Lipschitz constant of order $\Gamma_{n}=\frac{2d+1}{\sigma_{t}^{3}}+\frac{1}{2}$ which is only a mild restriction for $\sigma_{\min}\to 0$ as $n\to\infty$.
@@
    The network size in Theorem˜3.7 is not bounded from above. The empirical reference measure $\mathbb{P}^{\psi_{n}(Z)}$ in Theorem˜3.2 circumvents the need for concentration inequalities, which in turn avoids complexity bounds. Thus there is no trade-off in approximation power and network size.
  3. 3.
    Combining Theorem˜3.4 with Theorem˜3.2 and the proof strategy used in Theorem˜3.7 leads to a rate of convergence of $n^{-\frac{1+\alpha}{2\alpha+2+d}}$ for $p^{*}$ with unbounded support that satisfy a moment condition.
  4. 4.
    Compared to the literature on diffusion models and the first results on Flow Matching, we do not use early stopping times, but rather set $\sigma_{\min}>0$ according to $n.$ In terms of variance, both procedures are interchangeable: calculating $\sigma_{t_{*}}$ in (Fukumizu et al., 2025, Theorem 9) leads to exactly the same variance. However, early stopping leads to a bias induced by the mean function $\mu$, which is not $1$ at a time $t<1.$ This could be fixed using an adapted mean function, which is $1$ at the time of the early stopping, but this leads to much more complicated structures of $\mu.$
Theorem˜3.7 shows that Flow Matching achieves minimax optimal rates (up to logarithmic factors) for certain classes of unknown densities, giving some justification of their the empirical success. There are several limitations and potential avenues for further research in the results of this work.
First of all the networks used for Theorem˜3.7 are very large and do not correspond to the networks used in practice. While theoretical worst-case error bounds will always lead to networks that are not exactly small, the work of Yakovlev and Puchkin (2025) gives hope for improvement. Up to now it is an open question, whether a smaller network would come at the cost of a non-optimal rate. In addition, the size of the network and the overall result could benefit from the theoretical use of more sophisticated networks, like the U-net (Ronneberger et al., 2015) construction in Lipman et al. (2023).